In Thailand, a firm named SunFly is planning to monopolize its streaming services. The firm has made a ฿3 million investment in their streaming offerings. In addition, the firm has faced a marginal cost of ฿10 for each streaming service it offers. The table below displays SunFly's monthly streaming service pricing as well as the quantity demanded for each pricing. 


Price (฿):

  1. ฿ 130
  2. ฿ 120
  3. ฿ 110
  4. ฿ 100
  5. ฿ 90
  6. ฿ 80
  7. ฿ 70
  8. ฿ 60
  9. ฿ 50


Quantity demanded:

  1. 10,000
  2. 20,000
  3. 30,000
  4. 40,000
  5. 50,000
  6. 60,000
  7. 70,000
  8. 80,000
  9. 90,000


a.) Calculate total revenue, marginal revenue, total cost, and profit using the information given. What amount should a monopolistic business charge in order to maximize profits?


b.) Illustrate (draw) a graph for the marginal-revenue, marginal-cost, and demand curves based on your solution in (a). The marginal revenue and marginal-cost curves intersect at what price and quantity? Furthemore, give some explainations of your graph.


1
Expert's answer
2021-08-11T09:18:12-0400

Solution:

a.). TR = P "\times" Q

MR = "\frac{\partial TR} {\partial Q}"

TC = MC "\times" Q

Profit = TR – TC

The above has been calculated as per the below table:



 

 

A monopolistic business should charge a price of 70 in order to maximize profits.

This is the price where MR = MC, which is the profit-maximizing condition for a monopoly.

 

 

b.). The marginal revenue and marginal cost curves intersect at a price of 70 and quantity of 70000.

A monopolist will choose the profit-maximizing output where the MR intersects or equals the MC. A monopolist will then draw a straight vertical line from output to the demand curve and select the corresponding price, which is the profit-maximizing price.


This is displayed by the below graph: